# VLA-Cache: Towards Efficient Vision-Language-Action Model via Adaptive Token Caching in Robotic Manipulation

### v1

@@ 5.3 Results on Real Robot
Table 3 illustrates the performance of VLA-Cache in real-world robotic tasks. Among the four tasks, PickPot shows a slightly lower success rate than the baseline, whereas VLA-Cache exceeds the baseline on the other three tasks. The method also achieves considerable reductions in FLOPs and inference time. Overall, VLA-Cache improves the average success rate by 2.4%, an outcome that, although unexpected, is potentially explained by a more robust baseline model, which we trained extensively using additional data and more training steps than the LIBERO fine-tuning. With improved robustness as a foundation, VLA-Cache’s ability to prune or reuse redundant tokens may further enhance the model’s resilience, thus yielding higher success rates.
@@ 6 Conclusion
In this paper, we introduce VLA-Cache, a training-free method for VLA that selectively reuses static tokens while filtering out task-relevant ones, reducing redundant computation without sacrificing accuracy. Additionally, our layer-adaptive token reuse strategy improves model success rates by adjusting token reuse based on attention concentration. Extensive experiments on two VLA models, OpenVLA and CogAct, across two simulation environments, LIBERO and SIMPLER, demonstrate that VLA-Cache achieves a 1.7$\times$ speedup while maintaining performance.
Appendix A Appendix
A.1 Complexity Analysis Details
@@
A.2 VLA Cache Inference
Having identified both static tokens and task-relevant tokens as described in Section 4.1, we now detail how these tokens are reused in the VLA decoder via KV caching. Below, we first present a high-level overview of the procedure, then introduce several formulas to clarify the partial update of key-value (KV) representations under our selective token reuse strategy.
Overall Procedure.
At time step $t-1$, while predicting the action for frame $I_{t-1}$, the model produces the attention weights (used to evaluate task relevance) and the KV cache, $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$ for each layer $l$. When a new frame $I_{t}$ arrives, we compare it with $I_{t-1}$ (6) to obtain a set of static tokens sorted by similarity scores, and further filter out any high task-relevance tokens (9). Simultaneously, we determine a layer-specific reuse ratio $\alpha^{l}$  (12) based on the entropy of attention distributions at time $t-1$. This yields:
$\mathcal{P}_{\mathrm{final}}\;\text{and}\;\{\alpha^{l}\}_{l=1}^{\Omega},$ (21)
where $\mathcal{P}_{\mathrm{final}}$ denotes the final subset of tokens eligible for reuse, and $\Omega$ is the total number of decoder layers.
Selective Token Reuse.
During decoding at time $t$, each decoder layer $l$ reuses only the top-$k$ tokens (determined by $\alpha^{l}$ and similarity ranking) from $\mathcal{P}_{\mathrm{final}}$. Let $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$ be this chosen set of reusable tokens, with positions (IDs) $\{i\mid i\in\mathcal{P}_{\mathrm{reuse}}\}$. For tokens not in $\mathcal{P}_{\mathrm{reuse}}$, we recompute their key-value (KV) representations from scratch. Concretely, let $\mathbf{H}_{t}^{l}\in\mathbb{R}^{L\times D}$ be the hidden states at layer $l$ for the new frame $I_{t}$. The per-layer self-attention requires three components: queries $\mathbf{Q}_{t}^{l}$, keys $\mathbf{K}_{t}^{l}$, and values $\mathbf{V}_{t}^{l}$. We define:
$\mathbf{Q}_{t}^{l}(i)\;=\;W_{Q}^{l}\,\mathbf{H}_{t}^{l}(i),$ (22)
for all token indices $i\in\{1,\ldots,L\}$, where $W_{Q}^{l}\in\mathbb{R}^{D\times D}$ is the learnable projection of queries.
For $\mathbf{K}_{t}^{l}$ and $\mathbf{V}_{t}^{l}$, we partially reuse cached values from time $t-1$:
$\mathbf{K}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{K}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{K}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (23)
$\mathbf{V}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{V}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{V}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (24)
where $W_{K}^{l},W_{V}^{l}\in\mathbb{R}^{D\times D}$ are the key and value projection matrices, respectively, and $\mathbf{K}_{t-1}^{l}(i),\mathbf{V}_{t-1}^{l}(i)$ are the cached key-value vectors from the previous time step. This partial-update scheme allows us to skip computation for those tokens deemed both static and non-task-critical.
Finally, the self-attention output at layer $l$ is computed as:
$\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})\;=\;\mathrm{Softmax}\!\Bigl{(}\tfrac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\;\mathbf{V}_{t}^{l}.$ (25)
Only the newly computed tokens update $\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}$, while reused tokens retain their entries from $\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}$.
@@ A.2 VLA Cache Inference
Listing A.2 outlines our partial token reuse within the VLA decoder:
  1. 1.
    Position IDs and Causal Masks. We maintain a cache_position array reflecting the positions of tokens that require computation. Tokens not recomputed do not receive new position IDs, enabling a pruned causal mask aligned with the updated token set.
  2. 2.
    Computing Rotary Embeddings. Once the reusable tokens are omitted from re-encoding, we apply a rotary embedding function to the reduced set of token embeddings to generate position-dependent features for the next layer.
  3. 3.
    KV Cache Updates. Only newly computed tokens update their corresponding key-value entries in $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$. Reused tokens keep their entries from $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$. This partial-update mechanism leverages the permutation-invariance property of Transformers, ensuring consistent attention outputs even if certain tokens are omitted from recomputation.
Since Transformers naturally support skipping previously processed tokens in a self-attention cache, our layer-adaptive token reuse is readily integrated with standard caching strategies. Notably, the most significant reduction in computation occurs when generating the first action token at time $t$. After that, the autoregressive procedure continues as usual, incurring no additional overhead for subsequent tokens.
Efficiency and Effect on Performance.
By leveraging selective token reuse, VLA-Cache decreases the total number of tokens that must be re-encoded and re-attended at each time step $t$. This leads to a substantial reduction in decoding time, especially given the short sequences in robotic action prediction. Meanwhile, the complexity of predicting subsequent tokens in the same time step follows the usual KV-caching paradigm, adding minimal overhead. In practice, this method preserves task accuracy by forcing recomputation of highly dynamic or task-critical tokens while reusing the remainder. We present a detailed evaluation in Section 5, demonstrating that VLA-Cache achieves notable speedups with negligible performance degradation.
A.3 Finetuning Data for Real Robot Experiments
@@
The preprocessing procedure was conducted as follows: Initially, the recorded frames underwent center cropping, reducing the resolution from 1280×720 to 912×720, followed by resizing to 224×224. Subsequently, episodes were manually selected based on a visual inspection of the video sequences generated from the recorded frames. To minimize potential biases associated with excessively long episodes, those exceeding 250 steps were excluded from the dataset. Furthermore, steps in which all recorded action values were zero were removed to ensure data relevance and integrity.
A.4 Simulated Evaluation Detail.
LIBERO Task Definitions.
SIMPLER Task Definitions.
Implementation Details.
A.5 Additional Simulation Results
Result of Subtask on LIBERO Spatial Task Suit.
@@ A.5 Additional Simulation Results
task1 task2 task3 task4 task5 task6 task7 task8 task9 task10 Avg
Baseline (OpenVLA) 90 90 84 96 70 90 96 76 82 70 84.4 1.888 52.37
SparseVLM 88 60 90 90 60 82 90 92 72 74 79.8 1.367 88.08
FastV 92 90 90 94 58 92 90 80 78 70 83.4 1.888 54.00
VLA-Cache 90 90 88 94 66 84 94 84 76 72 83.8 1.382 32.22
@@
Figure 7 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache in LIBERO.
A.6 Real-World Evaluation Detail.
Task Definitions.
  1. 1.
    Pick Up Orange Pot (single-instruction): The robot's goal is to grasp the orange pot and lift it completely off the table. We collected 218 valid demonstrations for the training dataset, with slight random adjustments to the initial positions of the pot and the robotic arm in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  2. 2.
    Place Blue Cube in Box (single-instruction): The robot's goal is to place the held blue cube into the target container. We collected 212 valid demonstrations for the training dataset, with randomized container positions and robotic arm initial configurations in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  3. 3.
    Put Sausage in Blue Pan (single-instruction): The robot's goal is to stably place the held sausage into the blue pan. We collected 219 valid demonstrations for the training dataset, with randomized pan coordinates and robotic arm joint angles across trials. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  4. 4.
    Wipe Table (single-instruction): The robotic arm's goal is to sweep scattered items into a fixed-position dustpan using a broom. For the training dataset, we collected 187 valid demonstrations featuring randomized placements of simulated items (e.g., fries/cheese) on the table and varied initial joint configurations of the robotic arm, while the dustpan location remained fixed. This task explicitly incorporates dynamic environmental variations to rigorously test generalization capabilities. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
Visualization Results.
Figure 8 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache on the Kinova Jaco Robot Arm.
Figure 8: Comparison between baseline(OpenVLA) and VLA-Cache in real environment

### v2

@@ 5.4 Results on Real Robot
Table 5 illustrates the performance of VLA-Cache in real-world robotic tasks. Among the four tasks, PickPot shows a slightly lower success rate than the baseline, whereas VLA-Cache exceeds the baseline on other three tasks. The method also achieves considerable reductions in FLOPs and inference time. Overall, VLA-Cache improves the average success rate by 2.4%, likely due to reduced interference from redundant visual tokens and enhanced decision robustness. With improved robustness as a foundation, VLA-Cache’s ability to prune or reuse redundant tokens may further enhance the model’s resilience, thus yielding higher success rates.
Performance under Dynamic Background. To assess robustness, we introduced background motion (e.g., human hands and moving objects) in the PickPot task. As shown in Table 7, success rate of baseline dropped from 95% to 80% under noise. With VLA-Cache, the same success rate was maintained while reducing FLOPs by 42% and latency by 35%. These results highlight VLA-Cache’s ability to filter out transient or irrelevant tokens, preserving both efficiency and stability in dynamic real-world settings. Figure 4 visualizes this effect, showing robust token selection despite background disturbances.
@@ 6 Conclusion
[…]
Acknowledgments
This work was supported in part by the Australian Research Council under Projects DP240101848 and FT230100549. We would also like to thank the anonymous reviewers and area chair for their constructive feedback, which helped improve the clarity and quality of this paper.
Appendix A Limitations
In this section, we discuss the potential limitations of the proposed VLA-Cache: (i) In dynamic environments with substantial background or object motion, the number of non-reusable tokens increases, reducing acceleration gains. As visualized in Figure 4, dynamic regions (highlighted in yellow) require full recomputation. (ii) Our experiments focus on three state-of-the-art open-source VLA architectures (OpenVLA [11], CogAct [19], OpenVLA-OFT [20]) based on LLaMA2 [15] decoders. The applicability of VLA-Cache to emerging VLA systems with different backbones (e.g., Gemma2 [16] in $\pi_{0}$ [24]) or more complex VLA systems remains an open direction for future work.
Appendix B Impact Statement
While VLA-Cache demonstrates robust performance even under dynamic conditions, we emphasize the importance of careful deployment and continuous monitoring to ensure safe, interpretable, and reliable behavior in real-world robotic systems.
Appendix C Complexity Analysis Details
@@
Appendix D Inference Detail of VLA-Cache
Inference Procedure.
VLA-Cache accelerates robotic action prediction by reusing static visual tokens across frames during inference. At each timestep $t$, current visual tokens $\mathbf{H}_{t}$ are compared with those from the previous frame to identify unchanged regions. Tokens that are both visually static and task-irrelevant are reused via cached key-value (KV) entries, while task-relevant or dynamic tokens are recomputed. This selective reuse substantially reduces visual processing cost without altering model architecture or training.
Token Reuse Mechanism.
Our implementation modifies the VLA decoder’s forward pass as follows:
  • Position and Attention Masking. We maintain a cache_position array to mark tokens requiring recomputation. Static tokens retain their previous position encodings, allowing attention masks to be pruned to match the reduced token set.
  • Rotary Embedding. For recomputed tokens, rotary embeddings are applied to introduce positional information. Tokens that are skipped retain their previous encoded states.
  • Dynamic Cache Updates. Newly computed tokens update their respective $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$ entries, while reused tokens inherit values from the previous frame’s cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$. Due to the permutation invariance of Transformers, this partial update yields valid attention results.
This strategy is fully compatible with standard KV caching in autoregressive decoding. The largest computational gain occurs when generating the first action token at each timestep; subsequent tokens are decoded autoregressively without additional cost.
Experimental Settings.
All experiments are conducted using OpenVLA with 256 visual tokens. Unless specified otherwise, we use a static token similarity threshold $\tau=0.996$, top-$k=100$ for retained static tokens, and a task-relevance threshold $\tau_{\text{task}}=0.5$. These parameters are applied consistently across all simulated and real-world settings, including SIMPLER with CogAct. For real-world Jaco2 experiments, we slightly reduce the similarity threshold to $\tau=0.85$ to accommodate environmental noise. Training on the real robot used LoRA-based fine-tuning for 50,000 steps, and all evaluations were performed on an NVIDIA RTX 4090 GPU.
Appendix E More Experiment Results
E.1 Simulation Experiment
LIBERO Task Definitions.
SIMPLER Task Definitions.
@@ E.1 Simulation Experiment
E.2 Additional Simulation Results
Result of Subtask on LIBERO Spatial Task Suit.
Method Success Rate % $\uparrow$ FLOPs $\downarrow$ CUDA Time (ms) $\downarrow$
1 2 3 4 5 6 7 8 9 10 Avg
Baseline (OpenVLA) 90 90 84 96 70 90 96 76 82 70 84.4 1.888 52.37
SparseVLM 88 60 90 90 60 82 90 92 72 74 79.8 1.367 88.08
FastV 92 90 90 94 58 92 90 80 78 70 83.4 1.888 54.00
VLA-Cache 90 90 88 94 66 84 94 84 76 72 83.8 1.382 32.22
Table 7: Real-world PickPot task under dynamic background (human/object motion).
Method Success (%) $\uparrow$ FLOPs $\downarrow$ Latency (ms) $\downarrow$
Baseline (OpenVLA) 95 1.800 68.03
+ Noise 80 1.807 68.22
+ Noise, VLA-Cache 80 1.275 50.59
Visualization Results.
Figure 5 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache in LIBERO. More visualized results are available in the supplementary materials.
E.3 Additional Ablations and Comparisons
Attention vs. object–mask proxies for task relevance.
To validate whether attention score is a reliable proxy for task relevance, we compared our default attention–based filtering with an object-mask alternative using Efficient Track Anything[39]. On LIBERO-Spatial with OpenVLA-OFT, attention-based proxy yields both higher success rate and lower latency than the mask variant:
Method SR (%) $\uparrow$ FLOPs(T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
OpenVLA-OFT 97.8 3.99 78.35 65.44
   + VLA-Cache (Attention) 98.3 3.04 61.12 81.67
   + VLA-Cache (Object Mask) 87.4 3.15 87.49 64.78
Table 8: Attention vs. object-mask proxies for task relevance on LIBERO-Spatial.
While object masks provide spatial localization, they can miss fine-grained or contextual signals essential for manipulation, especially with background clutter or small, relevant parts. In contrast, attention scores are dynamically produced and tightly coupled with the model’s internal reasoning, providing a lightweight, task-adaptive proxy for relevance.
Sensitivity to static-token budget $k$ and relevance threshold $\tau$.
We further analyze the trade-off between accuracy and efficiency by varying the number of static tokens $k$ (with $\tau{=}0.5$ fixed) and the task-relevance threshold $\tau$ (with k=100k{=}100 fixed) on LIBERO-Spatial using OpenVLA-OFT. Results indicate robustness across a wide range; our default $(k{=}100,\ \tau{=}0.5)$ offers a strong balance.
Method SR (%) $\uparrow$ FLOPs (T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
OpenVLA-OFT 97.8 3.995 78.35 65.44
   + VLA-Cache (k=50k{=}50) 97.6 3.332 66.82 77.23
   + VLA-Cache (k=80k{=}80) 97.8 3.226 66.55 78.66
   + VLA-Cache (k=100k{=}100) 98.2 3.156 64.88 79.88
   + VLA-Cache (k=120k{=}120) 98.0 3.109 62.99 80.56
   + VLA-Cache (k=150k{=}150) 97.4 3.043 61.12 81.67
   + VLA-Cache (k=180k{=}180) 96.6 2.936 60.46 82.51
Table 9: Varying the static-token budget $k$ (with $\tau{=}0.5$).
Method SR (%) $\uparrow$ FLOPs (T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
OpenVLA-OFT 97.8 3.995 78.35 65.44
   + VLA-Cache ($\tau{=}0.2$) 95.6 3.384 66.93 76.74
   + VLA-Cache ($\tau{=}0.3$) 96.2 3.283 67.03 79.31
   + VLA-Cache ($\tau{=}0.4$) 98.0 3.204 66.38 79.79
   + VLA-Cache ($\tau{=}0.5$) 98.2 3.156 64.88 79.88
   + VLA-Cache ($\tau{=}0.6$) 98.6 3.131 64.27 81.98
   + VLA-Cache ($\tau{=}0.7$) 98.4 3.068 63.12 82.96
Table 10: Varying the relevance threshold $\tau$ (with k=100k{=}100).
Overall, efficiency (FLOPs and latency) improves monotonically with larger $k$ and $\tau$, while success rate remains consistently high, corroborating the stability of VLA-Cache across sensitivity settings. Our default $(k{=}100,\ \tau{=}0.5)$ is used for all main results unless stated otherwise.
Applicability to diffusion-based and alternative VLA architectures.
VLA-Cache is designed to accelerate the language-decoder stage of VLAs and is therefore not directly applicable to models without a vision–language backbone, such as standalone diffusion policies. However, for hybrid architectures that combine a VLM with a diffusion-based policy head (e.g., CogACT [19]), our method remains fully compatible and has shown consistent efficiency gains and stable task performance in the SIMPLER environment, as shown in Table 3. Since VLA-Cache operates by reusing temporally static visual tokens before action generation, it can be integrated with diffusion-based or transformer-style models (e.g., RDT, DiT) that exhibit inter-frame redundancy, offering a promising direction for future generalization.
E.4 Real Robot Experiment
@@
Task Definitions and Success Criteria.
We design four single-instruction tasks for the real-world robot setting to evaluate manipulation performance across diverse object types and interaction modes. The task definitions are as follows:
  1. 1.
    Pick Up Orange Pot (PickPot): The robot's goal is to grasp the orange pot and lift it completely off the table. A trial is successful if the pot is grasped and raised with visible clearance from the surface. We collected 218 valid demonstrations for the training dataset, with slight random adjustments to the initial positions of the pot and the robotic arm in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  2. 2.
    Place Blue Cube in Box (PlaceCube): The robot's goal is to place the held blue cube into the target container. A trial is successful if the cube is fully inside the container at the terminal step without contact-induced ejection. We collected 212 valid demonstrations for the training dataset, with randomized container positions and robotic arm initial configurations in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  3. 3.
    Put Sausage in Blue Pan (PutSausage): The robot's goal is to stably place the held sausage into the blue pan. Success requires the sausage to rest inside the pan boundary at the terminal step. We collected 219 valid demonstrations for the training dataset, with randomized pan coordinates and robotic arm joint angles across trials. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  4. 4.
    Wipe Table (WipeTable): The robotic arm's goal is to sweep scattered items into a fixed-position dustpan using a broom. Success requires all designated items to be inside the dustpan area at termination. For the training dataset, we collected 187 valid demonstrations featuring randomized placements of simulated items (e.g., fries/cheese) on the table and varied initial joint configurations of the robotic arm, while the dustpan location remained fixed. This task explicitly incorporates dynamic environmental variations to rigorously test generalization capabilities. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
Trial Protocol and Randomization.
We run 20 trials for PickPot and PutSausage, and 30 trials for PlaceCube and WipeTable, for a total of 100 trials per method. For each trial we randomize the initial robot configuration and the object placement within a bounded region of the workspace. Trials are single-attempt with a fixed horizon; there is no human intervention or reset within a trial. Outcomes are recorded as binary success or failure per the criteria above.
OpenVLA OpenVLA + VLA-Cache
Task Success Failure SR (%) Success Failure SR (%)
PickPot (20) 19 1 95.0 18 2 90.0
PlaceCube (30) 25 5 83.3 27 3 90.0
PutSausage (20) 16 4 80.0 17 3 85.0
WipeTable (30) 21 9 70.0 22 8 73.3
Total (100) 81 19 82.1 84 16 84.6
Table 11: Real-world results with trial counts and success rates.
Results with Counts and Rates.
Table 11 reports per-task successes and failures, along with the corresponding success rate. Average success is computed across all $100$ trials.
Figure 7: Comparison between baseline (OpenVLA) and VLA-Cache in real environment.
@@
Figure 7 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache on the Kinova Jaco Robot Arm. More visualized results are available in the supplementary materials.